Suppose a relative has promised to give you $1,000 as a wedding gift the day you get engaged. Assuming a constant interest rate of 6%, consider the present and future values of this gift, depending on when you become engaged. Complete the first row of the table by determining the value of the gift in one and two years if you become engaged today. Date Received - Present Value - Value in One Year - Value In Two Years Today 1,000 Fill in Fill in In 1 year Fill in 1,000 Fill in In 2 years Fill in Fill In 1,000 Complete the first column of the table by computing the present value of the gift if you get engaged in one year or two years. The present value of the gift is (greater, smaller) if you get engaged in one year than it is if you get engaged in two years.
Answer #1
Date Received Present Value($) Value in One Year($) Value in Two Year($)
Today 1000 1060 [i.e. 1000(F/P,6%,1) = 1000*1.060] 1124 [i.e. 1000(F/P,6%,2) = 1000*1.124]
In 1 year 943.4 [i.e. 1000(P/F,6%,1) = 1000*0.9434] 1000 1060.38 [i.e. 943.4(F/P,6%,2) = 1000*1.124
In 2 years 890 [i.e. 1000(P/F,6%,2) = 1000*0.8900] 943.4 [i.e. 890 (F/P,6%,1) = 1000*1.060] 1000